OpenStudy (anonymous):

how would i write a compound inequality and graph the time a cake must bake is beteween 25 minutes and 30 minutes,inclusive?

OpenStudy (mathstudent55):

Inclusive means that the time is from 25 to 30 minutes, including the 25 and 30 minutes. That means greater than or equal to and less than or equal to. Call the baking time t: \[t \ge 25 \space and \space t \le 30\] This can be written as \[25 \le t \le 30\]